Find the additive inverse of -9 on the 12-hour clock.
Found 3 solutions by stanbon, babygirlwalton, jim_thompson5910:
Answer by stanbon(75887)

Find the additive inverse of -9 on the 12-hour clock.
----
Ans: -3
Why?
-9+-3 = -12 = 0
==================
Chers,
Stan H.
===========================

Answer by babygirlwalton(3)

9 is correct answer i submitted it on my quiz -3 was wrong

Answer by jim_thompson5910(35256)

Add 12 to -9 to get 12+(-9) = 3

So -9 and 3 are the same number on the clock (ie going backwards 9 hours is the same as going forwards 3 hours...you land on 3 each time)

So you must find a number x such that

3 + x = 0

Solve for x:

3 + x = 0

3 + x - 3 = 0 - 3

x = -3

So -3 is the inverse of 3

Add 12 to -3 to get 12+(-3) = 9

This means that 9 is the inverse of -9
-------------------------------------------------------
Useful shortcut:

The additive inverse of 'x' is -x

Ex: x = 6 ---> additive inverse: -x = -6

Ex: x = 7 ---> additive inverse: -x = -7

etc etc